# LM-Searcher: Cross-domain Neural Architecture Search with LLMs via Unified Numerical Encoding

### v2

@@ 4.3 Out-of-Domain Tasks Experimental Details
Model Base Model Validation Test
LM-Searcher-1B LLaMA-3.1-1B 94.51 93.85
LM-Searcher-3B LLaMA-3.1-3B 94.60 93.85
@@
Table 5: Performance metrics for LM-Searcher with different base model size on CIFAR10.
@@ 5.2 Impacts of LLM Size
To investigate the impact of LLM size on performance, we evaluate LM-Searcher with different model scales, with results shown in Tab. 5. We train models of different sizes and evaluate the results on NAS-Bench-101. Our results show a consistent improvement as the LLM size increases, indicating that larger models can better capture the promising architecture patterns.
@@ 5.3 Architecture-performance Mapping
To further validate LM-Searcher’s capability in identifying efficient architectures through pattern analysis of historical experimental data, we conducted an ablation study where we intentionally randomized the architecture-performance mapping in NAS-Bench-201. This manipulation introduces noise into the information provided to the LLM. As depicted in Tab. 6, the performance degrades significantly when the architecture-performance pairs provided to LM-Searcher are corrupted by noise. This result shows that LM-Searcher is able to utilize Ncode-performance pairs provided in the context to effectively search architectures. More experimental analysis are provied in Appendix B.
@@ Appendix A In-Domain Tasks Experimental Details
To benchmark LM-Searcher’s performance against other NAS approaches. We utilize the widely-used NAS-Bench-201 Dong and Yang (2020). To mitigate the effects of randomness, each search experiment is conducted over five independent runs. As depicted in Table B.2, LM-Searcher achieves competitive performance across all the dataset.
Appendix B Experimental Analysis
B.1 LM-Searcher Search behavior
@@
B.2 LM-Searcher Attention Pattern

### v3

@@ 3.2 LM-Searcher
CIFAR10 CIFAR100 ImageNet Kvasir ISIC 2017 Multi-Concept Customization IWSLT’14 De-En NB-ASR
@@ 4.3 Out-of-Domain Tasks Experimental Details
Model Base Model NAS-Bench-101 CIFAR100 ImageNet
LM-Searcher-1B LLaMA-3.1-1B 93.85 72.31 73.52
LM-Searcher-3B LLaMA-3.1-3B 93.85 72.60 73.93
@@
Table 5: Performance metrics for LM-Searcher with different base model size.
@@ 5.2 Impacts of LLM Size
To investigate the impact of LLM size on performance, we evaluate LM-Searcher with different model scales, with results shown in Tab. 5. We train models at different scales and report test accuracy on NAS-Bench-101, CIFAR-100, and ImageNet. Our results show a consistent improvement as the LLM size increases, indicating that larger models can better capture the promising architecture patterns.
@@ 5.3 Architecture-performance Mapping
To further validate LM-Searcher’s capability in identifying efficient architectures through pattern analysis of historical experimental data, we conducted an ablation study where we intentionally randomized the architecture-performance mapping in NAS-Bench-201. This manipulation introduces noise into the information provided to the LLM. As depicted in Tab. 6, the performance degrades significantly when the architecture-performance pairs provided to LM-Searcher are corrupted by noise. This result shows that LM-Searcher is able to utilize Ncode-performance pairs provided in the context to effectively search architectures. More experimental analysis are provied in Appendix C.
5.4 Trade-off Between Specialization and Generalization
We conducted in-domain fine-tuning with LM-Searcher by encoding architectures with more domain-specific information rather than simple digit encoding. For example, a structure originally encoded as 333123 is now encoded as |conv3X3~0|+|conv3X3~0|+|conv3X3~1|+|skip_connect~0|+|conv1X1~1|+|conv3X3~2|, where the numbers following the tilde (~) indicate which previous feature (node) the operator (edge in the DAG) connects to. As shown in Tab. 8, after fine-tuning on 1000 samples created using this encoding method, LM-Searcher achieves performance that approaches or even surpasses existing state-of-the-art methods. However, finetuning the model with such domain-specific samples restricts cross-domain generalization, resulting in performance degradation in other domains (0.97% decline on IWSLT’14 De-En machine translation and 0.13% decline on NAS-Bench-ASR).
@@ 5.4 Trade-off Between Specialization and Generalization
[…]
DARTS (Liu et al., 2018a) 54.30 39.77 16.32 - -
SGNAS (Huang and Chu, 2021) 93.53 70.31 44.98 - -
DiNAS (Asthana et al., 2024) 94.37 73.51 45.51 - -
AG-Net (Lukasik et al., 2022) 94.37 73.51 46.42 - -
LM-Searcher-FT 94.36 73.51 46.54 33.05 21.57
[…]
Table 8: Performance on in-domain image classification (CIFAR-10, CIFAR-100, ImageNet-16 from NAS-Bench-201) and out-of-domain tasks (IWSLT’14 De–En machine translation, NB-ASR speech recognition). LM-Searcher-FT is fine-tuned on 1,000 domain-specific samples.
@@ Appendix A In-Domain Tasks Experimental Details
To benchmark LM-Searcher’s performance against other NAS approaches. We utilize the widely-used NAS-Bench-201 Dong and Yang (2020). To mitigate the effects of randomness, each search experiment is conducted over five independent runs. As depicted in Table C.2, LM-Searcher achieves competitive performance across all the dataset.
Appendix B Metric Explanation
S-measure ($S_{\alpha}$)
S-measure evaluates the structural similarity between the predicted segmentation and ground truth. It considers both region-aware ($S_{r}$) and object-aware ($S_{o}$) structural similarities:
$S_{\alpha}=\alpha\cdot S_{o}+(1-\alpha)\cdot S_{r}$
where $\alpha$ is typically set to 0.5. The object-aware similarity focuses on the foreground object consistency, while region-aware similarity evaluates the overall structural information.
E-measure ($E_{\phi}$)
The Enhanced-alignment measure captures both pixel-level matching and image-level statistics. It is computed as:
$E_{\phi}=\frac{1}{W\times H}\sum_{x=1}^{W}\sum_{y=1}^{H}\phi_{FM}(x,y)$
where $\phi_{FM}$ is the enhanced alignment matrix, and $W$, $H$ are the width and height of the image.
Jaccard Index (Jac)
Jaccard Index is also known as Intersection over Union (IoU), it measures the overlap between predicted and ground truth regions:
$\text{Jaccard}=\frac{|P\cap G|}{|P\cup G|}$
where $P$ is the predicted segmentation and $G$ is the ground truth.
F-measure ($F_{\beta}^{\omega}$)
The weighted F-measure combines precision and recall with adaptive weights:
$F_{\beta}^{\omega}=\frac{(1+\beta^{2})\cdot\text{Precision}^{\omega}\cdot\text{Recall}^{\omega}}{\beta^{2}\cdot\text{Precision}^{\omega}+\text{Recall}^{\omega}}$
where the superscript $\omega$ indicates weighted versions that emphasize errors in important regions.
Boundary Error Rate (BER)
The boundary error rate specifically evaluates the accuracy of boundary detection, calculated as:
$\text{BER}=\frac{|B_{P}\setminus B_{G}|+|B_{G}\setminus B_{P}|}{|B_{G}|}$
where $B_{P}$ and $B_{G}$ are the predicted and ground truth boundaries, respectively.
Bilingual Evaluation Understudy (BLEU)
Bilingual Evaluation Understudy (BLEU) is an automatic metric for evaluating machine translation and other text generation tasks by measuring how closely a system’s output matches one or more reference translations. It computes modified n-gram precision: counts of overlapping 1- to 4-grams between the candidate and references, clipped so repeated n-grams aren’t over-rewarded. These precisions are combined via a geometric mean and multiplied by a brevity penalty that downweights outputs shorter than the references.
Phoneme Error Rate (PER).
Phoneme Error Rate (PER) is a standard metric for evaluating phoneme-level recognition or alignment systems. Given a reference phoneme sequence and a hypothesized sequence, PER is computed as the edit distance between the two sequences—decomposed into substitutions (S), deletions (D), and insertions (I)—normalized by the length of the reference sequence (N). Formally,
$\mathrm{PER}=\frac{S+D+I}{N}\times 100\%,$
where $S$, $D$, and $I$ are obtained via an optimal alignment (e.g., dynamic programming for Levenshtein distance).
Appendix C Experimental Analysis
C.1 LM-Searcher Search behavior
@@
C.2 LM-Searcher Attention Pattern